OpenStudy (anonymous):

I need help solving for x in this equation: 2(x+8)=6x-14 Please and thank you! I need help solving for x in this equation: 2(x+8)=6x-14 Please and thank you! @Mathematics

OpenStudy (anonymous):

2(x+8)=6x-14 x+8=3x-7 divide by 2 x-3x=-7-8 subtract 3x; subtract 8 -2x=-15 simplify x=15/2 divide by -2

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

np :})

hero (hero):

np = nope

OpenStudy (saifoo.khan):

np = not possible.

OpenStudy (anonymous):

NP refers to "nondeterministic polynomial time"

hero (hero):

NP = Not Probable, Not Possible, or Not Plausible

OpenStudy (anonymous):

2(x+8)=6x-14 2x+16=6x-14 16+14=6x-2x 30=4x then divide by 4 and you get

OpenStudy (anonymous):

7.5 Thanks Also!! :)